OpenStudy (anonymous):

5(whole number)1/5 ÷ 6(whole number)1/6 =t÷7(whole number)1/7

OpenStudy (anonymous):

\[5\frac{ 1 }{5 }\times 7\frac{ 1 }{7 }= 6\frac{ 1 }{6 }t \] I want to check my answer! Please try this out!

OpenStudy (anonymous):

5\frac{ 1 }{5 }\times 7\frac{ 1 }{7 }= 6\frac{ 1 }{6 }t I want to check my answer! Please try this out!